# Continuous Temporal Domain Generalization

### v1

@@ A.1.4 Complexity Analysis
In our framework, coordinate transformations between spaces are implemented using an autoencoder achieved by MLP. The complexity of these transformations is described by a linear relationship in terms of $N$, expressed as $\mathcal{O}(2(Nd+E)+F)$, where $N$ denotes the number of parameters in the predictive models (i.e., $\theta$), $d$ denotes the number of neurons in the final MLP layer, $E$ denotes for the number of parameters remaining in the autoencoder, and $F$ denotes the number of parameters in the Koopman operator. Note that in many deep learning tasks, significant portions of the model’s layers function as representation learning and can be shared, resulting in a small $N$.
Table 2: Ablation Test Results for Different Datasets
@@ A.1.5 Ablation Study
As shown in the table, each component of constraints contributes to robust performance across all tested datasets. The removal of particular elements has marked effects on model stability and accuracy. Specifically, the absence of $L_{pred2}$ and $L_{dyna}$ constraints leads to significant performance degradation, highlighting their essential roles in learning the correct dynamics of the model through continuous temporal domain data. Removing other constraints leads to unpredictable results on certain datasets or a decline in effectiveness, indicating their importance in ensuring system effectiveness and stability under various input conditions. The considerable increase in error observed with the ✗K.S on the 2-Moons dataset demonstrates the challenges associated with modeling nonlinear dynamics directly in the parameter space, and the out-of-memory (OOM) problems with Rot-MNIST and Cyclone upon the removal of the Koopman Space suggest that this component is not only pivotal but also computationally demanding.
Figure 7: Scalability Analysis w.r.t Number of Parameters and Number of Domains
@@ A.1.6 Scalability Analysis
A.1.7 Limitations
Our approach focuses on the gradual concept drift, which may require some additional technical tools in the face of sudden change.
A.2 Supplementary Figure for Analysis and Control Experiments
Figure 8: Predictive Model Trajectories of the Baseline Model; Left: It is clear that DRAIN as a probabilistic model cannot capture the continuous dynamics, which is presented as jump from one state to another; Right: DeepODE presents some continuity (red interpolated trajectories), but the dynamics are not correct at all.
Figure 9: Extremely Long-Term Extrapolated Predictive Model Trajectories in Uncontrolled and Controlled Settings; Left: $\mathcal{K}$ as learnable; Right: $\mathcal{K}=B-B^{T}$ with $B$ as learnable
@@ A.2 Supplementary Figure for Analysis and Control Experiments
Appendix B Experiments on Discrete Temporal Domains
@@
In TDG experiments, we follow the problem definition in [1]. Formally, a sequence of data domains is collected over discrete time intervals. Each domain, denoted as $\mathcal{D}_{t}$, corresponds to a dataset collected at a specific time $t$ where $t=1,2,...,T$. We have a sequence of domains $\{\mathcal{D}_{1},\mathcal{D}_{2},...,\mathcal{D}_{T}\}$, where each domain $\mathcal{D}_{t}=\{(x_{i}^{t},y_{i}^{t})\}_{i=1}^{N_{t}}$ consists of $N_{t}$ instances with features $x_{i}^{t}\in X_{t}$ and labels $y_{i}^{t}\in Y_{t}$, and $X_{t}$ and $Y_{t}$ represent the random variables for features and targets. The goal in TDG is to train a predictive model $g(X_{t};\theta_{t})$ on these historical domains, and modeling the dynamics of $\{\theta_{1},\theta_{2},...,\theta_{T}\}$, then predict the $\theta_{T+1}$ for the unseen future domain $\mathcal{D}_{T+1}$ at time $t=T+1$, finally do the prediction task $Y_{T+1}^{*}=g(X_{T+1};\theta_{T+1})$.
We have replicated the experimental setups typically used in TDG research. Specifically, we used datasets, settings, and hyperparameters from the DRAIN [1]. We compare with the following classification datasets: Rotated Moons (2-Moons), Rotated MNIST (Rot-MNIST), Online News Popularity (ONP), and Shuttle; and the following regression datasets: House prices dataset (House), Appliances energy prediction dataset (Appliance). The first two datasets differ from the data used in our CTDG main experiments in Exp. 5, as here they are time-regularly distributed and only have one 1-step future domain to predict. We keep the architecture of the predictive model for each dataset the same as DRAIN [1].
@@
The results are shown in Table 3. The Koodos model consistently outperforms all baselines across various datasets, demonstrating its robustness and superior adaptability also made to traditional temporal domain generalization tasks. Notably, Koodos significantly reduces error margins, especially in real-world datasets such as House and Appliance. The exceptional performance of Koodos can be attributed to its foundational continuous dynamic system design, which effectively captures and synchronizes the dynamics of data and model parameters. which is efficient and applicable in both discrete and continuous temporal domains.
@@ Appendix C Theorem Proofs
Assumption 2.
We make the following assumptions:
1. Continuous TDG Model (CTDG): The system’s state $x(t)$ evolves according to the ODE:
$\frac{dx(t)}{dt}=f(x(t),t),$ (16)
where $f$ is a smooth function.
2. Discrete TDG Model (DTDG): The system’s state $h_{t}$ is updated at discrete time steps $t$ according to:
$h_{t+1}=\phi(h_{t},x_{t}),$ (17)
where $\phi$ is a smooth function and $x_{t}$ is the input at time $t$.
3. Error Analysis: Let $\Delta t$ be the time step for the DTDG updates. The continuous-time method approximates the state at time $t+\Delta t$ using the integral of $f(x(t),t)$ over the interval $[t,t+\Delta t]$. The discrete-time method approximates the state at time $t+\Delta t$ using a step-wise update without considering the integral of changes within the interval.
Intuitively speaking, the CDTG and DTDG models can be parameterized by e.g., Neural ODE and RNN, respectively.
Theorem 3.
Given the assumptions, a continuous-time method provides a more accurate approximation of the temporal distribution drift than a discrete-time method, in terms of the error accumulated over time intervals.
@@ 1. Continuous TDG Model (CTDG): The system’s state $x(t)$ evolves according to the ODE:
Below is our formal proof of our theorem.
For CTDG method, the state $x(t+\Delta t)$ can be approximated by the ODE integral:
@@
Using the Taylor series expansion for $f(x(t),t)$, we get:
$x(t+\Delta t)=x(t)+f(x(t),t)\Delta t+\mathcal{O}(\Delta t^{2}).$ (19)
The error in this approximation is:
$\text{Error}_{\text{CTDG}}=\mathcal{O}(\Delta t^{2}).$ (20)
For DTDG method, the state $h_{t+1}$ is updated as:
$h_{t+1}=\phi(h_{t},x_{t}).$ (21)
Assuming $\phi$ approximates $f$ over discrete intervals, we can write:
$h_{t+1}=h_{t}+\phi(h_{t},x_{t})\Delta t+\mathcal{O}(\Delta t^{2}).$ (22)
To analyze the error, let’s approximate the update rule in a continuous form. Assume $\phi$ is a smooth function that approximates the continuous-time dynamics $f$.
Using Taylor series expansion around $t$ with step size $\Delta t$, we get:
$h_{t+1}=h_{t}+\Delta t\cdot\phi(h_{t},x_{t})+\frac{(\Delta t)^{2}}{2!}\cdot\frac{\partial\phi}{\partial t}+\mathcal{O}((\Delta t)^{3}).$ (23)
For simplicity, consider the first-order approximation:
$h_{t+1}\approx h_{t}+\Delta t\cdot\phi(h_{t},x_{t}).$ (24)
The difference between the true continuous state $x(t+\Delta t)$ and the discrete state $h_{t+1}$ is:
@@
Expanding $f(x(t),t)$ using Taylor series, we get:
$f(x(\tau),\tau)\approx f(x(t),t)+(\tau-t)\frac{\partial f}{\partial t}+\mathcal{O}((\tau-t)^{2}).$ (26)
The integral of $f(x(\tau),\tau)$ over $[t,t+\Delta t]$ is:
[…]
Therefore:
$x(t+\Delta t)\approx x(t)+\Delta t\cdot f(x(t),t)+\frac{(\Delta t)^{2}}{2}\cdot\frac{\partial f}{\partial t}+\mathcal{O}((\Delta t)^{3}).$ (28)
Comparing this with the DTDG update:
$h_{t+1}\approx h_{t}+\Delta t\cdot\phi(h_{t},x_{t}).$ (29)
The error term $\epsilon_{\text{RNN}}$ is:
$\epsilon_{\text{RNN}}=\left(x(t)+\Delta t\cdot f(x(t),t)+\frac{(\Delta t)^{2}}{2}\cdot\frac{\partial f}{\partial t}+\mathcal{O}((\Delta t)^{3})\right)-\left(h_{t}+\Delta t\cdot\phi(h_{t},x_{t})\right).$ (30)
Assuming $h_{t}\approx x(t)$ and $\phi(h_{t},x_{t})\approx f(x(t),t)$:
$\epsilon_{\text{RNN}}\approx\frac{(\Delta t)^{2}}{2}\cdot\frac{\partial f}{\partial t}+\mathcal{O}((\Delta t)^{3}).$ (31)
The dominant error term is $\mathcal{O}(\Delta t)$ from the first-order approximation, making the DTDG error primarily first-order:
$\text{Error}_{\text{DTDG}}=\mathcal{O}(\Delta t).$ (32)
For the CDTG method, the error term is quadratic in $\Delta t$:
$\text{Error}_{\text{CTDG}}=\mathcal{O}(\Delta t^{2}).$ (33)
For the DTDG method, the error term is linear in $\Delta t$:
$\text{Error}_{\text{DTDG}}=\mathcal{O}(\Delta t).$ (34)
Since $\Delta t^{2}\leq\Delta t$ for sufficiently small $\Delta t$, we have:
$\text{Error}_{\text{CTDG}}\leq\text{Error}_{\text{DTDG}}.$ (35)
Thus, under the given assumptions, the continuous-time method using an ODE has a smaller or equal error compared to the discrete-time method using an RNN for approximating temporal distribution drift. This demonstrates that the ODE method provides a more accurate approximation due to its consideration of the integral of changes over time, reducing the accumulation of errors compared to the step-wise updates of the RNN.

### v2

@@ A.1.4 Complexity analysis
In our framework, coordinate transformations between spaces are implemented using an autoencoder achieved by MLP. The complexity of these transformations is described by a linear relationship in terms of $N$, expressed as $\mathcal{O}(2(Nd+E)+F)$, where $N$ denotes the number of parameters $\theta$ in the predictive models, $d$ denotes the number of neurons in the first MLP layer, $E$ denotes for the number of parameters remaining in the autoencoder, and $F$ denotes the number of parameters in the Koopman operator. In many deep learning tasks, significant portions of the model’s layers function as representation learning and can be shared, resulting in a small $N$ and controlled model complexity.
Table 2: Ablation test results for different datasets.
@@ A.1.5 Ablation study
As shown in the table, each component of constraints contributes to robust performance across all tested datasets. The removal of particular elements has marked effects on model stability and accuracy. Specifically, the absence of $L_{integ}$ and $L_{dyna}$ constraints leads to significant performance degradation, highlighting their essential roles in learning the correct dynamics of the model through continuous temporal domain data. Removing other constraints leads to unpredictable results on certain datasets or a decline in effectiveness, indicating their importance in ensuring system effectiveness and stability under various input conditions. The considerable increase in error observed with the ✗K.S on the 2-Moons dataset demonstrates the challenges associated with modeling nonlinear dynamics directly in the parameter space, and the out-of-memory (OOM) problems with Rot-MNIST and Cyclone upon the removal of the Koopman Space suggest that this component is not only pivotal but also computationally demanding.
Figure 8: Scalability analysis w.r.t the number of parameters and the number of domains.
Table 3: Cost of training and testing time.
Model Train Time (s) Test Time (s)
DRAIN 67 <1
DeepODE 289 1
CIDA 610 3.05
TKNets 948 2
Koodos 566 2.83
@@ A.1.6 Scalability analysis
We further conduct the model scalability analysis by comparing the running time of Koodos with other state-of-the-art baselines: DRAIN, DeepODE, CIDA, and TKNets in the 2-Moons dataset. As shown in Table 3, our model strikes a good balance between training time and effectiveness.
Figure 9: Sensitivity analysis.
A.1.7 Sensitivity analysis
We conducted the sensitivity analysis on 2-Moons and House datasets to understand the hyperparameters in Koodos: loss weights $(\alpha,\beta,\gamma)$ and the dimensions $n$ of the Koopman operator $\mathcal{K}$. The loss weights are set based on the magnitude of each loss term. For instance, in the 2-Moons dataset, the cross-entropy loss $L_{intri}$ and $L_{integ}$ are approximately 1 after convergence, the $L_{recon}$ and $L_{consis}$ in the Model space are around 0.01, and $L_{dyna}$ in the Koopman space is around 0.1. Accordingly, we set the initial values of $\alpha,\beta,\gamma$ to 1, 100, and 10, respectively. We then adjust each term independently within its respective range: $\alpha$ and $\gamma$ are varied from 1 to 100, and $\beta$ from 10 to 1000. The dimension $n$ of the Koopman operator varies from 16 to 2048. Fig. 9 shows the results.
It can be seen that: (1) Koodos exhibits stable behavior with controlled variance across a wide range of hyperparameter values, evidencing its insensitivity to hyperparameter variations. (2) Setting the loss weights of $(\alpha,\beta,\gamma)$ by the magnitude of each loss term is sufficient for achieving good performance, and fine-tuning the weights may give better results. (3) A hundreds-dimension approximation of the Koopman operator is sufficient to achieve good results. Too high may lead to overfitting.
A.1.8 Limitations
The limitation arises from the assumptions inherent in the research area of the temporal domain generalization framework. TDG assumes that domain pattern drifts follow certain predictable, smooth patterns, allowing for modeling future changes as a sequence. CTDG extends TDG to continuous time, further generalizing the problem to handle domains collected at arbitrary times. Thus, their assumptions are aligned in focusing on smooth, predictable concept drift. To address other kinds of drift, alternative frameworks may be required. However, different frameworks can be combined to address the full spectrum of domain generalization challenges. Exploring such a comprehensive approach may represent a promising research direction.
@@ A.1.8 Limitations
Appendix B Experiments on discrete temporal domain generalization
@@
In TDG experiments, we follow the problem definition in DRAIN [2]. Formally, a sequence of data domains is collected over discrete time intervals. Each domain, denoted as $\mathcal{D}_{t}$, corresponds to a dataset collected at a specific time $t$ where $t=1,2,...,T$. We have a sequence of domains $\{\mathcal{D}_{1},\mathcal{D}_{2},...,\mathcal{D}_{T}\}$, where each domain $\mathcal{D}_{t}=\{(x_{i}^{t},y_{i}^{t})\}_{i=1}^{N_{t}}$ consists of $N_{t}$ instances with features $x_{i}^{t}\in X_{t}$ and labels $y_{i}^{t}\in Y_{t}$, and $X_{t}$ and $Y_{t}$ represent the random variables for features and targets. The goal in TDG is to train a predictive model $g(X_{t};\theta_{t})$ on these historical domains, and modeling the dynamics of $\{\theta_{1},\theta_{2},...,\theta_{T}\}$, then predict the $\theta_{T+1}$ for the unseen future domain $\mathcal{D}_{T+1}$ at time $t=T+1$, finally do the prediction task $Y_{T+1}^{*}=g(X_{T+1};\theta_{T+1})$.
We have replicated the experimental setups typically used in TDG research. Specifically, we used datasets, settings, and hyperparameters from the DRAIN [2]. We compare with the following classification datasets: Rotated Moons (2-Moons), Rotated MNIST (Rot-MNIST), Online News Popularity (ONP), and Shuttle; and the following regression datasets: House prices dataset (House), Appliances energy prediction dataset (Appliance). Note that the first two datasets differ from datasets used in our CTDG main experiments in Exp. 5, as here they are time-regularly distributed and only have 1-step future domain to predict. We keep the architecture of the predictive model for each dataset the same as DRAIN [2].
@@
The results are shown in Table 4. The Koodos model consistently outperforms all baselines across various datasets, demonstrating its robustness and superior adaptability also made to traditional temporal domain generalization tasks. Notably, Koodos significantly reduces error margins, especially in real-world datasets such as House and Appliance. The exceptional performance of Koodos can be attributed to its fundamental continuous dynamic system design, which effectively captures and synchronizes dynamic changes in data and models.
@@ Appendix C Theorem proofs
Assumptions
1. Continuous Data Distribution Drift: - Let $p(t)$ be the data distribution at time $t$. - Assume $p(t)$ is continuously differentiable, and the drift $\frac{dp(t)}{dt}$ is Lipschitz continuous with constant $L$.
2. Training and Target Domains: - The model is trained on a sequence of training domains $\{p(t_{i})\}_{i=1}^{T}$ where $t_{1}<t_{2}<\ldots<t_{T}$. - The target domain is at a future time $s$, with the data distribution $p(s)$. - The time intervals $\Delta t_{i}=t_{i+1}-t_{i}$ can be arbitrary.
3. Model Training: - For the ODE model, let $x(t)$ be the state at time $t$, governed by the ODE:
$\frac{dx(t)}{dt}=f(x(t),t),$
where $f$ is a smooth function. - For the RNN model, let $h_{t_{i}}$ be the state at discrete time $t_{i}$, updated by:
$h_{t_{i+1}}=h_{t_{i}}+\Delta t_{i}\cdot\phi(h_{t_{i}},x_{t_{i}}),$
where $\phi$ is a smooth function and $x_{t_{i}}$ is the input at time $t_{i}$.
4. Generalization Error: - The generalization error on the target domain $p(s)$ depends on the accumulated training errors and the error propagation from the last training domain $p(t_{T})$ to the target domain $p(s)$.
Theorem 3 (Formal version of Theorem 2).
Given the assumptions, a continuous-time method using an Ordinary Differential Equation (ODE) provides a smaller or equal generalization error on an unseen target domain compared to a discrete-time method using a Recurrent Neural Network (RNN), due to its more accurate approximation of the data distribution drift over time.
@@ 3. Model Training: - For the ODE model, let $x(t)$ be the state at time $t$, governed by the ODE:
Below is the formal proof:
Accumulated Training Error
1. ODE Model: - The state $x(t+\Delta t_{i})$ can be approximated by the ODE integral:
@@
- Using the Taylor series expansion for $f(x(t),t)$:
$x(t_{i}+\Delta t_{i})=x(t_{i})+f(x(t_{i}),t_{i})\Delta t_{i}+\frac{(\Delta t_{i})^{2}}{2}\frac{\partial f}{\partial t}+\mathcal{O}((\Delta t_{i})^{3}).$ (17)
- The error for one step is:
$\text{Error}_{\text{ODE, step}}=\mathcal{O}((\Delta t_{i})^{2}).$ (18)
- The accumulated error over $T$ steps is:
$\text{Error}_{\text{ODE, train}}=\sum_{i=1}^{T-1}\mathcal{O}((\Delta t_{i})^{2}).$ (19)
- Given that the intervals $\Delta t_{i}$ can be large, the accumulated error becomes:
$\text{Error}_{\text{ODE, train}}=\mathcal{O}\left(\sum_{i=1}^{T-1}(\Delta t_{i})^{2}\right).$ (20)
2. RNN Model: - The state $h_{t_{i+1}}$ is updated as:
$h_{t_{i+1}}=h_{t_{i}}+\Delta t_{i}\cdot\phi(h_{t_{i}},x_{t_{i}}).$ (21)
- Assuming $h_{t_{i}}\approx x(t_{i})$ and $\phi(h_{t_{i}},x_{t_{i}})\approx f(x(t_{i}),t_{i})$, the error for one step is:
$\text{Error}_{\text{RNN, step}}=\mathcal{O}(\Delta t_{i}).$ (22)
- The accumulated error over $T$ steps is:
$\text{Error}_{\text{RNN, train}}=\sum_{i=1}^{T-1}\mathcal{O}(\Delta t_{i}).$ (23)
- Given that the intervals $\Delta t_{i}$ can be large, the accumulated error becomes:
$\text{Error}_{\text{RNN, train}}=\mathcal{O}\left(\sum_{i=1}^{T-1}\Delta t_{i}\right).$ (24)
Error Propagation to Target Domain
1. ODE Model: - To handle an arbitrary future time $s$, we integrate the error over many small steps from $t_{T}$ to $s$:
@@ 2. RNN Model: - The state $h_{t_{i+1}}$ is updated as:
- Since $\tau$ ranges from $t_{T}$ to $s$:
$\text{Error}_{\text{ODE, target}}\approx\text{Error}_{\text{ODE, train}}+\mathcal{O}((s-t_{T})^{3}).$ (26)
2. RNN Model: - Similarly, the error propagation to the target domain $s$ involves many small steps:
$\text{Error}_{\text{RNN, target}}\approx\text{Error}_{\text{RNN, train}}+\sum_{k=0}^{K-1}\mathcal{O}(\Delta t_{k}),$ (27)
- where $K$ is the number of small steps from $t_{T}$ to $s$ and $\Delta t_{k}$ are the small intervals:
$\text{Error}_{\text{RNN, target}}\approx\text{Error}_{\text{RNN, train}}+\mathcal{O}(s-t_{T}).$ (28)
Generalization Error Comparison
- For the ODE model:
$\text{Error}_{\text{ODE, target}}\approx\mathcal{O}\left(\sum_{i=1}^{T-1}(\Delta t_{i})^{2}\right)+\mathcal{O}((s-t_{T})^{3}).$ (29)
- For the RNN model:
$\text{Error}_{\text{RNN, target}}\approx\mathcal{O}\left(\sum_{i=1}^{T-1}\Delta t_{i}\right)+\mathcal{O}(s-t_{T}).$ (30)
- Since $(\Delta t_{i})^{2}$ is much smaller than $\Delta t_{i}$ for any $\Delta t_{i}$, and $(s-t_{T})^{3}$ is much smaller than $s-t_{T}$, we have:
$\text{Error}_{\text{ODE, target}}\leq\text{Error}_{\text{RNN, target}}.$ (31)